OpenStudy (anonymous):

Evaluate 30 ÷ (6 - 1) • 2

OpenStudy (angela210793):

30/5*2=30/10=3

OpenStudy (anonymous):

You have to follow bodmas (so division first)-  = 6x2 =12. (i think.)